OpenStudy (anonymous):

A uniform steel wire X and Y, suspended at the same horizontal level, are connected by a light and cross-sectional area A, A weight of 30 N is placed at O, on the road OX : OY = 1 : 2. Both wires are stretched and the rod XY then remains horizontal. If he wire X has a young modulus E1 of 1.0 * 10^11 pa, calculate the yound modulus of E2 of the wire Y assuming the elastic limit is not exceeded for both wires.

OpenStudy (anonymous):



OpenStudy (anonymous):

@Hero @Preetha @ghazi @AravindG

OpenStudy (aravindg):

I think I saw a similiar question in NCERT text chapter 11 under topic of Young's modulus and all..take a look

OpenStudy (anonymous):

can u give me the link of the book?

OpenStudy (aravindg):

you are not from India?

OpenStudy (anonymous):

nope

OpenStudy (aravindg):

i suggest you google it..its easy to get NCERT CLASS 11 PHYSICS PART 2

OpenStudy (anonymous):

do they have the solutions

OpenStudy (aravindg):

yes i think .....

OpenStudy (aravindg):

worth checking

OpenStudy (aravindg):

infact the book is very nice for reference

OpenStudy (anonymous):

can u change the language?

OpenStudy (aravindg):

ofcourse eng is below scroll down

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

the question isnt there

OpenStudy (anonymous):

Ok.. so since the xy is horizontal.. it means after stretching, both the bars have exactly the same length.. so i can say \[l1+\Delta l1 = l2 + \Delta l2 \] next step we find the stress that is coming on each of them.. that should be easy.. since the ration of ox: oy is given to us..

OpenStudy (anonymous):

can u show the working please:)

OpenStudy (anonymous):

try doing it :P.. first you tell me how much will be the stress that comes on X and Y ?

OpenStudy (anonymous):

stress = force/ area how do i distrbte the force

OpenStudy (anonymous):

first find the force.. force equals the weight that each one would experience.. the weight is at a ratio of 2:1 .. hence the distribution will be in inverse ratio right? 1:2.. cause if weight is far away, its effect is smaller!

OpenStudy (anonymous):

hence Fx : Fy = 2:1

OpenStudy (anonymous):

so now calculate Fx and Fy

OpenStudy (anonymous):

Fx = 2/3 Fy Fy = 1/3 Fx ?

OpenStudy (anonymous):

total force is known to you.. so find the ratios now!

OpenStudy (anonymous):

Fx = 2/3 Ftotal :P and Fy = 1/3 Ftotal

OpenStudy (anonymous):

Fx = 20N and Fy = 10N

OpenStudy (anonymous):

ok.. :D.. next.. find the stress on X and since young's modulus of X is given.. find the strain.. Similarly do the same thing for Y.. but young's modulus is not given.. divide the two equations to get rid of A

OpenStudy (anonymous):

thanks alot Strain = F/A*E1 strain in X = 20/A* 1.0 * 10^11 strain is equal so E2 = stress/ strain E2 = 10/A *A*1.0*10^11/20 E2 = 5 * 10^11 pa

OpenStudy (anonymous):

woah wait.. how did you get the strain for Y ?

OpenStudy (anonymous):

arent they equal

OpenStudy (anonymous):

nope :-/.. why would the strains be equal?

OpenStudy (anonymous):

same mterial

OpenStudy (anonymous):

now you have to use the first equatoin.. that i put.. but use lx and ly notation instead of l1 and l2 notation

OpenStudy (anonymous):

they are NOT SAME MATERIAL.. IF THEY WERE.. THEY WOULD HAVE SAME YOUNG"S MODULUS :P

OpenStudy (anonymous):

i dont undrstnd >< cn u shw the working

OpenStudy (anonymous):

use the first equation .. try doing it.. see you need to get rid of \[\Delta ly/ ly\] in that equation i asked you to build by dividing the two equations

OpenStudy (anonymous):

so here are the two equations \[Fx / Fy = Ex \Delta x ly / Ey \Delta ylx \] and \[lx + \Delta x = ly + \Delta y\] somehow use them

OpenStudy (anonymous):

ok il use thnks fr ur help :)

OpenStudy (anonymous):

No problimo amigo :D